LDG Commissioning: Performance Verification - 24 April 2010 UT

Night 6:

Observer: JHeidt (LBTO), PHibon (LBTO), JHill (LBTO)
Telescope Operator: JHill (LBTO)
Telescope Support: JUrban (LBTO)
Instrument Support: JMorris (LBTO)
Adaptive Secondary Operator: MXompero (LBTO)

Summary

Lost to weather - humidity, clouds and snow showers all night.

Preparations

JHoward made more Elevation pad adjustments to HBS.

DAshby rebalanced after some catwalk steel was added to the telescope.

Jay and Kevin were working on a failed fan in the utility bldg chiller.

22:49 Stop snow melt after it has been running all day. Snow melt plus a few patches of sunshine have melted the 5 cm of snow from the roof. However, there could be more snow approaching from the North. The snow/rime is cleaned off of 3 sides of the building. Jeff reports a few pieces of ice still clinging to the lightning ball.

Details

01:20 Open rear vent doors manually halfway to cool chamber (presently +5) down toward ambient (presently -7) Building is parked at AZ=285, so there are nice photos of the shadow of LBT on High Peak.

01:42 Start Mirror Ventilation. Glass temp is +5 degC.

01:50 Auto-raise SX primary after panic

02:00 Sky is partly cloudy. It is NOT snowing. We are presently closed for humidity, and the primary mirrors being 12 degC above outside ambient. T=-7.3 degC, D=-7.3 degC. Wind from North gusting to 15 m/sec. This wind is also kicking up powdery snow from the forest floor.

02:10 Clear the LDG chain CCW limit.

02:15 Home the AGw4 probe.

02:20 Authorize IRTC direct on Left side.

02:30 Pascale and Jochen send a few test presets to verify that everything seems to be working.

03:00 Close rear vent doors as some snow flurries have started. Mirror Ventilation is still running for the moment.

04:45 Stop Mirror Ventilation as glass is down to +1 degC. Presently we are closed by humidity, clouds, winds gusting to 20 m/sec, and blowing snow (perhaps recycled out of the forest).

06:27 Set HBThermalLoopOffset back to the nominal value of -10 degC. Stop and Restart ECS. James left us with a "k" value of 130 (compared to the previous 170) which is providing 122.3 Bar. This helps the oil temperature situation, as do the colder temperatures when we have less oil flow in the bearings.

07:27 Reset encoders at AZ=120 EL=90.

07:30 Wind has moderated (mostly below 10 m/sec). Temperature has increased to -5.9 degC. But we still have thick low clouds and 100% humidity.

09:00 It has started snowing again, so it looks like we are done for the night.

09:40 Retracted SX M2 and deployed LBC-Blue for Andrew's daytime tests on Saturday. Authorize LBC on both sides.

-- JohnHill - 23 Apr 2010
Topic revision: r7 - 24 Apr 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback